# On the Convergence of Multi-objective Optimization under Generalized Smoothness

### v3

@@ 5.2 Stochastic setting
Let Assumptions 1, 2 and 3 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}}$, $\alpha\sim\mathcal{O}(\epsilon^{9})$, $\beta\sim\mathcal{O}(\epsilon^{4})$, $\rho\sim\mathcal{O}(\epsilon^{2})$, $n_{s}\sim\mathcal{O}(\epsilon^{-6}),N\sim\mathcal{O}(\epsilon^{-2}),$ and $T\sim\Theta(\epsilon^{-11})$, and all the parameters satisfy the requirements in Theorem 2. We then have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\sim\mathcal{O}(\epsilon),$ with the probability at least $1-\delta$.
@@ 6 Experiments
In this experiment, we evaluate the performance of the Cityscapes dataset [10], which involves 2 pixel-wise tasks: 7-class semantic segmentation (Task 1) and depth estimation (Task 2). Following the same experiment setup of [32], we build a SegNet [2] as the model, comparing the performance of MGDA [12], PCGrad [34], GradDrop [8], CAGrad [24], MoCo [13], MoDo [6], Nash-MTL [27], SDMGrad [32] with our methods, GSMGrad and GSMGrad-FA with the warm-start initialization. We utilize the metric $\mathbf{\Delta m\%}$ to reflect the overall performance, which considers the average per-task performance drop versus the single-task (STL) baseline to assess methods. It can be observed in Figure 2 that GSMGrad has a better result in task 2 and a much more balanced performance. Meanwhile, the proposed GSMGrad-FA is much faster than GSMGrad as shown in Table 2 in the Appendix.
In addition, we also illustrate the relationship between the gradient norm and the local smoothness for each task. To do so, we compute them according to the method displayed in Section H.3 in [35]. We scatter the local smoothness constant against gradient norms in Figure 2 for the semantic segmentation task and depth estimation task in Figure 3 (in the appendix), respectively. Both results demonstrate a positive correlation between them, which further substantiates the necessity of our analysis. More experimental details can be found in Appendix A.
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ STL 74.01 93.16 0.0125 27.77 MGDA [12] 68.84 91.54 0.0309 33.50 44.14 PCGrad [34] 75.13 93.48 0.0154 42.07 18.29 GradDrop [8] 75.27 93.53 0.0157 47.54 23.73 CAGrad [24] 75.16 93.48 0.0141 37.60 11.64 MoCo [13] 75.42 93.55 0.0149 34.19 9.90 MoDo [6] 74.55 93.32 0.0159 41.51 18.89 Nash-MTL [27] 75.41 93.66 0.0129 35.02 6.82 SDMGrad [32] 74.53 93.52 0.0137 34.01 7.79 GSMGrad 75.41 93.46 0.0133 31.07 3.93 GSMGrad-FA 74.38 93.24 0.0160 41.78 19.44
@@ 7 Conclusion
In this paper, we investigate the multi-objective problem with a more challenging, relaxed and realistic $\ell$-smooth assumption. We propose the first efficient MOO algorithm GSMGrad and its stochastic variant SGSMGrad for this problem. We provide the convergence guarantee for both algorithms to find an $\epsilon$-accurate Pareto stationary point with $\epsilon$-level average/iteration-wise CA distance. Extensive experiments are conducted to validate our theoretical results.
@@ A.1 Relation between gradient norms and the local smoothness
A.2 Running time comparison between GSMGrad and GSMGrad-FA
We compare the average running time of the proposed algorithms, GSMGrad and GSMGrad-FA. The time in Table 2 is an average of the total running time over epochs (in minutes). The result solidifies the advantage of the fast approximation.
Method Average running time
GSMGrad 2.93
GSMGrad-FA 1.93
Table 2: Average running time comparison between GSMGrad and GSMGrad-FA.
@@ A.3 Implementation details
Multi-task learning on Cityscapes dataset. Following the experiment setup in [32], we train our method for 200 epochs, using SGD optimizers for both model parameters and weights, and the batch size for Cityscapes is 8. We compute the averaged test performance over the last 10 epochs as the final performance measure. We fix the $\beta=0.5$ and do a grid search on hyperparameters including $N\in[10,20,40,50],\alpha\in[0.0001,0.0002,0.0005,0.001]$, and $\rho\in[0.01,0.05,0.1,0.2,0.5,0.6,0.7,0.8,0.9,1]$ and choose the best result from them. It turns out our best performance of GSMGrad is based on the choice that $N=40,\alpha=0.0005,\beta=0.5$, and $\rho=0.5$. The choice of hyperparameters for GSMGrad-FA turns out to be the same as that for GSMGrad. All experiments are run on NVIDIA RTX A6000.
@@ Appendix B Algorithm
We show our GSMGrad with Fast Approximation (GSMGrad-FA):
Algorithm 4 GSMGrad with Fast Approximation (GSMGrad-FA)

### v4

@@ 5.2 Stochastic setting
Let Assumptions 1, 2 and 3 hold. Set $\beta^{\prime}\leq\frac{1}{M^{2}}$, $\alpha\leq\mathcal{O}(\epsilon^{9})$, $\beta\leq\mathcal{O}(\epsilon^{4})$, $\rho=\mathcal{O}(\epsilon^{2})$, $n_{s}\geq\mathcal{O}(\epsilon^{-6}),N\geq\mathcal{O}(\epsilon^{-2}),$ and $T\geq\Theta(\epsilon^{-11})$, and all the parameters satisfy the requirements in Theorem 2. We then have $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\leq\mathcal{O}(\epsilon),$ with the probability at least $1-\delta$.
@@
Method Segmentation Depth Surface Normal MR $\downarrow$ $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 STL 38.30 63.76 0.6754 0.2780 25.01 19.21 30.14 57.20 69.15 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 7.89 5.59 SI 38.45 64.27 0.5354 0.2201 27.60 23.37 22.53 48.57 62.32 7.33 4.39 RLW [23] 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 9.89 7.78 DWA [26] 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 7.11 3.57 UW [20] 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 7.11 4.05 MGDA [12] 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 5.56 1.38 MoCo [13] 40.30 66.07 0.5575 0.2135 26.67 21.83 25.61 51.78 64.85 5.00 0.16 MoDo [6] 35.28 62.62 0.5821 0.2405 25.65 20.33 28.04 54.86 67.37 7.55 0.49 Nash-MTL [29] 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 3.33 -4.04 FAMO [24] 38.88 64.90 0.5474 0.2194 25.06 19.57 29.21 56.61 68.98 3.22 -4.10 MGDA-warm start 40.57 67.17 0.5240 0.2281 25.21 19.74 28.74 55.79 68.21 2.78 -4.42
Table 2: Multi-task supervised learning on NYU-v2 dataset for different MOO methods comparison.
@@ 6 Experiments
In this experiment, we evaluate the performance of the Cityscapes[10] and NYU-v2 [33] datasets. The former involves 2 pixel-wise tasks: 7-class semantic segmentation (Task 1) and depth estimation (Task 2) while the latter involves 3 pixel-wise tasks: 13-class semantic segmentation, depth estimation and surface normal estimation. Following the same experiment setup of [35], we build a SegNet [2] as the model. We compare the performance of MGDA with a warm start, which is used to ensure a small CA distance per iteration and mitigate gradient conflicts, against popular MGDA-type methods including MGDA [12], PCGrad [37], GradDrop [8], CAGrad [25], MoCo [13], MoDo [6], Nash-MTL [29], and FAMO [24]. Since SDMGrad [35] incorporates a preference-based regularization within the MGDA framework, we exclude it from the tables to ensure a fair comparison. We utilize the metric $\mathbf{\Delta m\%}$ to reflect the overall performance, which considers the average per-task performance drop versus the single-task (STL) baseline to assess methods. It can be observed in Table 2 and Figure 2 that MGDA-warm start has a much more balanced performance. Meanwhile, the proposed MGDA-FA is much faster as shown in Figure 4 in the Appendix.
We also illustrate the relationship between the gradient norm and the local smoothness for each task of the Cityscapes dataset. To do this, we compute them according to the method provided in Section H.3 in [38]. We scatter the local smoothness constant against gradient norms in Figure 1 for the semantic segmentation task and depth estimation task in Figure 3 (in the appendix), respectively. Both results demonstrate a positive correlation between them, which further substantiates the necessity of our analysis. More experimental details can be found in Appendix A.
Method Segmentation Depth $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ STL 74.01 93.16 0.0125 27.77 MGDA [12] 68.84 91.54 0.0309 33.50 44.14 PCGrad [37] 75.13 93.48 0.0154 42.07 18.29 GradDrop [8] 75.27 93.53 0.0157 47.54 23.73 CAGrad [25] 75.16 93.48 0.0141 37.60 11.64 MoCo [13] 75.42 93.55 0.0149 34.19 9.90 MoDo [6] 74.55 93.32 0.0159 41.51 18.89 Nash-MTL [29] 75.41 93.66 0.0129 35.02 6.82 FAMO [24] 74.54 93.29 0.0145 32.59 8.13 MGDA-warm start 75.41 93.46 0.0133 31.07 3.93
@@ 7 Conclusion
Building upon our observations of MGDA convergence under the $(L_{0},L_{1})$ smoothness condition, this paper provides a rigorous convergence analysis for both the fundamental MGDA and its stochastic variant under a more challenging, relaxed, and practical generalized $\ell$-smoothness assumption. Furthermore, we introduce a warm start progress to provide a more precise control over the iteration-wise CA distance. Our analysis also shows that an efficient variant named MGDA-FA, which uses only $\mathcal{O}(1)$ time and space, achieves the same performance guarantee as MGDA. We anticipate that the convergence analysis developed in this work will provide valuable insights for analyzing other MOO algorithms such as CAGrad, PCGrad, FairGrad and FAMO under the generalized smoothness condition. The warm-start strategy may be of independent interest to other single-loop MOO algorithms to achieve a sufficiently small iteration-wise CA distance.
@@ A.1 Relation between gradient norms and the local smoothness
A.2 Results and running time comparison between MGDA-warm start and MGDA-FA
We compare the results and average running time of the proposed algorithms, MGDA-warm start and MGDA-FA of the Cityscapes [10]. The time in Figure 4 is an average of the total running time over epochs (in minutes). The result solidifies the advantage of the fast approximation.
Method Segmentation Depth $\Delta m\%\downarrow$ Average running time mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ STL 74.01 93.16 0.0125 27.77 MGDA-warm start 75.41 93.46 0.0133 31.07 3.93 2.93 MGDA-FA 74.38 93.24 0.0160 41.78 19.44 1.93
Figure 4: Multi-task learning on Cityscapes dataset.
@@ A.3 Implementation details
Multi-task learning on Cityscapes dataset. Following the experiment setup in [35], we train our method for 200 epochs, using SGD optimizers for both model parameters and weights, and the batch size for Cityscapes is 8. We compute the averaged test performance over the last 10 epochs as the final performance measure. We fix the $\beta=0.5$ and do a grid search on hyperparameters including $N\in[10,20,40,50],\alpha\in[0.0001,0.0002,0.0005,0.001]$, and $\rho\in[0.01,0.05,0.1,0.2,0.5,0.6,0.7,0.8,0.9,1]$ and choose the best result from them. It turns out our best performance is based on the choice that $N=40,\alpha=0.0005,\beta=0.5$, and $\rho=0.5$. The choice of hyperparameters for MGDA-FA turns out to be the same as that for MGDA-warm start. All experiments are run on NVIDIA RTX A6000.
@@ Appendix B Algorithm
Algorithm 4 MGDA with Fast Approximation (MGDA-FA)